Starting at age 25, you deposit $2000 a year into an IRA account for retirement. Treat the yearly deposits into the account aplease show all work and steps thank you :)
Answer #1

Effective interest rate = e^0.05 - 1

= 1.051271096 - 1

= 0.051271096

a.

FW of IRA account = 2000*(F/A,5.1271096%,40)

= 2000* ((1 + 0.051271096)^40-1)/ 0.051271096

= 2000* ((1.051271096)^40-1)/ 0.051271096

= 249226.43 ~ 249226 (Nearest Dollar)

b.

FW of interest = 249226 - 2000*40 = 169226 (Nearest Dollar)